OpenStudy (anonymous):

how would i simplify -a+3 over -2a+6 using -1 as a factor

OpenStudy (anonymous):

\[\frac{-a+3}{-2a+6} = \frac{1(-a+3)}{2(-a+3)}\] I don't see where or why you'd want to factor out a -1 there.

OpenStudy (anonymous):

factor out -a+2

OpenStudy (anonymous):

ope i meant +3